40 likes | 170 Vues
This meeting report outlines the current status and preparations for the RHIC Run 11 as of January 4, 2011. The AGS is expected to achieve 60-65% polarization with 1.5*10^11 intensity. Critical ramp systems have been successfully tested, with smooth switching anticipated between 9/28 MHz cavities. Preparations for both downstream and upstream polarimeter tests are underway. The schedule includes a cryo cooldown period with minimal beam activity, leading to an optimistic start for beam setup on Friday night. Meetings for progress updates are scheduled for Friday morning and afternoon.
E N D
RHIC Status of pp Run11 Haixin Huang Time Meeting 01/04/2011
Critical Systems • AGS is robust to deliver 60-65% polarization with 1.5*1011 • AtR: beam has been dumped into W dump. • Ramp System • Tested successfully. • Two ramps in hands: one regular one and one with D’ shifted down • RHIC Injection • Switching between 9/28MHz cavities is expected to be smooth • RHIC CNI polarimeters • Dry run for down stream polarimeter has finished. Test run for upstream polarimeter will happen tomorrow. • Tune/orbit feedback on the ramp • Tune feedback is successful, orbit feedback will require beam for further test. The feedback system is expected to be ready for the first ramp.
Schedule • Jan. 3-9 Cryo cool down, not much beam activity is expected, just a few owl shifts over the weekend. • The most optimistic estimate is to start beam setup in blue Friday night. Power supply work will take two shifts per day for the weekend, probably also first a few days next week. • The 9AM meeting will start on Friday. • The 3:30PM meeting for beam development will start when we have beam time during day time.